Given a pool of 30 year fully-amortizing FRMs making monthly payments to investors with the following characteristics: WAC: 6% Pass-through rate: 5.5% Prepayment assumption: 200 PSA Loans were not seasoned before entering pool This MBS has been active for a few years in collecting payments from borrowers and making payments to investors. It's currently month 26 in the pool.(t=26) Starting pool balance month in month 26: 72,534,232 What is the starting pool balance for month 27 of this security? (Hint: N=360 when t=1, N=359 when t=2, N=358 when t=3....etc)
Answer #1

Starting pool balance month in month 26: 72,534,232

26th month payment = 72,534,232/[(1-1/((1+(0.06/12))^335))/(0.06/12)] =$446689.69

where 0.06=WAC, 335=(30*12-25) months

interest =(0.055/12)* 72,534,232 = $332448.56

Scheduled principal payment = 446689.69 - [(0.06/12)*72,534,232] =$84018.53 , where 0.06=WAC

Using 200% PSA model,

CPR = 2.00 *(26/30)*.06 =0.104 ,here 0.06*(26/30) when t=26 is used as per 100%PSA model standard

SMM= 1-[(1-.104)^(1/12)] =.0091095

So, prepaid principal = .0091095 *(72,534,232 - 84018) =$659985.22

Now, starting pool balance for month 27 of this security=72,534,232 - 84018.53 - 659985.22 = $71,790,228